Jesse made 26 out of 110 free throws. What percent of the free throws did Jesse make?
Answer by stanbon(75887)

Jesse made 26 out of 110 free throws. What percent of the free throws did Jesse make?
------
Solve:
x/100 = 26/110
-----
x = 2600/110
-----
x = 23.64%
=================
Cheers,
Stan H.